HYDROGEN ABSTRACTION FROM CH_{3}D BY CHLORINE RADICALS WITH VARYING KINECTIC EN-ERGY DISTRIBUTIONS

ANDREW E. BERKE, \underline{ETHAN H. VOLPA}, F. FLEMING CRIM, Chemistry Department, University of Wisconsin - Madison, Madison, Wisconsin 53706.

Vibrationally driven hydrogen abstraction from methane and its isotopologues has become nearly ubiquitous as a standard in gas phase reaction dynamics studies. To obtain a more complete view of the reaction dynamics of the hydrogen abstraction reaction of CH_{3}D with Chlorine radicals, we have studied how a dramatic change in the kinetic energy distribution of the Cl radicals affects the abstraction dynamics of our system. To this end, we have performed pumpprobe experiments, exciting the 2ν_{4} CH_{3}D antisymmetric C-H stretch overtone, photolyzing Cl_{2} with laser pulses of 309 nm, 355 nm, and 416 nm, then probing the products with (2+1) REMPI. We find both similarities and marked differences in the abstraction dynamics at the three different photolysis energies.